OpenStudy (anonymous):

how to solve for d d/4.2 = (d+30)/6

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

As for this question all you need to do is just cross multiply and then solve it!! :D \[\frac{ d }{ 4.2 } = \frac{ d+30 }{ 6 }\] \[6d = 4.2d + 126\] Now bring the variables to one side and then solve for d here!! :D \[d(6-4.2) = 126\] \[d = 126/1.8 \] Understood? :)